# PanoDreamer: 3D Panorama Synthesis from a Single Image

### v1

@@ 4.1 Single-Image Panorama Generation
During this stage, we fix $J_{T-1},\dots,J_{0}$ and find the optimal $L$ that minimizes Eq.4. $L$ influences both the first term, as the diffusion model is conditioned on it (see Eq. 6), and appears in the second term. The optimal solution to the first term is $L^{*}=L$. The idea is that if $L$ was used to produce the current $J_{T-1},\dots,J_{0}$, it is likely the best option for reproducing the same results. On the other hand, since the second term is quadratic, the solution is simply $L^{*}=J_{0}$. Although obtaining the solution to each term is straightforward, computing the optimal solution considering both terms is difficult. However, assuming that $L$ and $J_{0}$ are close to each other—i.e., MultiConDiffusion does not diverge significantly from the condition image in one pass—it is reasonable to assume that $L^{*}=J_{0}$ is close to the optimal solution for both terms.
@@
We perform the optimization by first initializing $J_{T}$ with Gaussian noise and $L$ by progressively inpainting the image. We then alternate between stages 1 and 2 iteratively until convergence. At the end of this process, we can use either $\tilde{J}_{0}$ or $\tilde{L}$ as the final result. Fig. 2 compares MultiConDiffusion with MultiDiffusion and progressive inpainting, while Fig. 3 shows the results of our approach across different iterations of the optimization process.
4.1.1 Panorama Generation Details
@@
We slightly modify the MultiConDiffusion process to adapt it for generating panoramas from a single image. Our goal is to produce a cylindrical panorama, so in this case, MultiConDiffusion operates in the cylindrical domain, and the sequence $J_{T},\dots,J_{0}$ is defined within this domain. Since the base diffusion model operates on perspective images, $F_{i}$ performs both cropping and projection from the cylindrical to the perspective domain. Similarly, $F_{i}^{-1}$ projects the pixels from the perspective image back to the cylindrical image, placing them in the appropriate locations.
@@
We experimented with bilinear interpolation during the projection; however, interpolation smoothed out the noise, which negatively affected the performance of the diffusion model. Therefore, we instead use nearest neighbor interpolation for both $F_{i}$ and $F_{i}^{-1}$. Additionally, we use an FOV of 45° for the perspective camera and carefully set the resolution of the cylindrical image to ensure a rough one-to-one mapping between the pixels of the cylindrical and perspective images.
@@
Figure 4: We compare the result of our method, PanoDepthFusion, against applying Depth Anything V2 (DA V2) [32] on the full image. The results obtained by DA V2 lacks details and is geometrically inconsistent. Our approach, on the other hand, produces highly detailed and consistent depth maps.
Figure 5: On the top left, we show the result of averaging the patch depth estimates. As seen, since the depth maps are relative, the depth from different patches are not consistent, producing results with clear edges. Since we initialize $G_{\theta_{i}}$ with identity line, patchwise average results are indeed our initial depth estimate during the optimization of Eq. 7. We also show our results after one, two, and four iterations of optimization. After only four iterations, the seems disappear.
4.2 Panorama Depth Estimation
Given the panoramic image $\tilde{J}_{0}$, our goal is to estimate its depth $D$. In recent years, several powerful monocular depth estimation methods [32, 31] have been introduced. These approaches can estimate highly detailed relative depth but typically perform best at a specific image size. Beyond this optimal resolution, they often produce results that lack detail and geometric consistency. Consequently, applying these methods directly to panorama depth estimation leads to poor results, as shown in Fig. 4.
We address this problem by obtaining $D$ through a combination of estimated depth maps on patches using an existing technique, $\Psi$, i.e., $\Psi(F_{i}(\tilde{J}_{0}))$. However, naïvely combining the patches (e.g., through averaging) leads to unsatisfactory results (see Fig. 5), as the patch depth estimates are relative and can be inconsistent. To overcome this, we pose the problem of obtaining panoramic depth from patch depth estimates as an optimization task. Our key insight is that the panoramic depth $D_{J}$ should be close to the estimated depth after it has been globally aligned through a parametric function. This can be formally written as:
$D^{*},\bm{\theta}^{*}=\operatorname*{arg\,min}_{D,\bm{\theta}}\sum_{i=1}^{n}\|F_{i}(D)-G_{\theta_{i}}(F_{i}(\tilde{J}_{0}))\|^{2},$ (7)
@@ 4.2 Panorama Depth Estimation
Solving for both $D$ and $\bm{\theta}$ simultaneously is challenging. Therefore, we propose performing this optimization through alternating minimization, consisting of two stages. In the first stage, we fix $\bm{\theta}$ and find the optimal $D$. Since the objective is quadratic, the solution can be obtained in closed form, similar to Eq. 3. The only difference is that $\Phi$, the diffusion model, is replaced with $\Psi$, and $W_{i}=\bm{1}$. In the second stage, we fix $D$ and find the optimal $\bm{\theta}$. This is a least-squares regression problem, which can be solved using standard packages.
@@
Figure 6: We compare the panoramas generated by MultiConDiffusion with those from other methods. Other approaches often result in sharp discontinuities and contextual inconsistencies. For instance, in the top example, the MultiDiffusion result shows a mismatch between the generated sky and the input sky.
Figure 7: We show the results of our approach on the same input image across multiple runs. As shown, our approach produces diverse yet consistent results.
Table 1: Numerical comparison of MultiConDiffusion against other panorama generation approaches. “Q-IQA” denotes CLIP-IQA+ quality, “Q-Align” denotes Q-Align quality, “A-Align” denotes Q-Align aesthetic score, “A-Clip” denotes CLIP aesthetic score, “C-Clip” denotes CLIP consistency, and “C-Style” denotes style consistency.
Method Q-IQA $\uparrow$ Q-Align $\uparrow$ A-Align $\uparrow$ A-Clip $\uparrow$ C-Clip $\uparrow$ C-Style (×$\times 10^{-3}$) $\downarrow$
Progressive 0.520 4.164 3.314 5.779 0.896 0.881
MultiDiffusion [1] 0.523 4.39 3.516 5.953 0.900 1.23
SyncDiffusion [13] 0.535 4.290 3.429 5.893 0.898 0.798
MultiConDiffusion (ours) 0.530 4.481 3.696 5.992 0.902 0.793
Table 2: Evaluation on rendered 3D scene. The evaluation metrics of 3D scene are the same as panorama evaluation.
Method Q-IQA $\uparrow$ Q-Align $\uparrow$ A-Align $\uparrow$ A-Clip $\uparrow$ C-Clip $\uparrow$ C-Style $\downarrow$
LucidDreamer 0.495 2.911 2.705 5.253 0.848 0.058
WonderJourney [36] 0.504 3.506 2.834 5.368 0.820 0.058
PanoDreamer (ours) 0.443 3.305 2.772 5.673 0.869 0.025
4.3 Inpainting and Optimization
@@
The estimated depth can now be used to project the cylindrical image into 3D space. However, when the scene is viewed from any position other than the panorama’s center of projection, occluded regions become visible. To address this, we utilize the layered depth image (LDI) inpainting approach by Shih et al. [22], which performs effective depth-aware texture inpainting while also providing the corresponding depth. We use a four-layer LDI representation (foreground, background, and two intermediate layers) based on agglomerative clustering by disparity.
@@
We then use these cylindrical layered images and depth maps to initialize a set of 3D Gaussians. Specifically, we assign a Gaussian to each pixel of the image at each layer and project them into 3D according to the corresponding depth. The color of each Gaussian is initialized based on the pixel color (without using spherical harmonics); we initialize the rotation matrix with identity, assign the scale following Paliwal et al. [20], and set the opacity to 0.5. During this process, we keep track of which Gaussians correspond to which layer, as this information is required for optimization.
@@
Next, we perform 3DGS optimization for 1000 iterations. To do this, we set up 240 evenly rotation cameras from the center of projection and project the layered images and depth maps to these cameras. During the optimization, we randomly sample one of these cameras and optimize the Gaussians according to their corresponding layer. Additionally, we composite all the four layers and use the composited image as a reference to optimize all the Gaussians. We use the original 3DGS reconstruction loss along with an $L_{2}$ loss between the rendered and layered depth maps. In addition, to be able to produce consistent results from novel views, we use the depth-based novel view loss, proposed by Zhu et al. [43].
Given this optimized 3DGS representation, we can synthesize novel views of the scene and produce coherent and seamless results.
5 Results
In this section, we compare our approach against state-of-the-art panorama generation and 3D scene generation methods, both visually and numerically. We demonstrate that our method generates more coherent panoramas and 3D scenes than other approaches. For evaluation, we compile a test set of 28 real and synthetic scenes sourced from LucidDreamer [3] and WonderJourney [36].
Figure 8: We compare renderings of PanoDreamer with LucidDreamer [3] and WonderJourney [36]. The input image and camera are at the leftmost column. For each methods, we render 3D scene from two novel views (orange and green cameras). As is shown in the figure, LucidDreamer and WonderJourney results are inconsistent, suffering severe seam-artifacts for novel views. In comparison, PanoDreamer is capable of generating coherent renderings from novel views. For more visual results and video comparison, please refer to our supplementary materials.
@@ 5 Results
For numerical evaluation, we employ the following three metrics: (1) Quality metrics. We assess the quality of generated panorama and 3D scenes using CLIP-IQA+ [27] and Q-Align [29] score. CLIP-IQA+ and Q-Align are built upon Contrastive Language-Image Pre-training (CLIP) [21] and large multi-modality models (LMMs) for image quality assessment, respectively. (2) Consistency metrics. To assess the consistency of the generated panoramas and 3D scene, we use CLIP and style metrics, where we compute the CLIP embedding similarity and style loss [7] between random pairs of non-overlapping window images in the panorama and 3D scene. (3) Aesthetic score. Similar to WonderWorld[35], we also evaluate the aesthetic score using CLIP aesthetic score and Q-Align aesthetic score.
5.1 Panorama Reconstruction Comparisons
In Fig. 6, we show visual results of MultiConDiffusion, and compare our method against SyncDiffusion [22], MultiDiffusion [1] and vanilla progressive inpainting using DMs. Note that, since these images are not as wide as cylindrical panoramas, we perform the optimization for 15 iterations instead of 20. As seen, progressive inpainting generates panoramic images with noticeable seams and strong inconsistency. The discontinuity improves in MultiDiffusion and SyncDiffusion, but they still struggle to preserve good contextual consistency across the entire panoramas. In contrast, MultiConDiffusion can generate coherent and seamless panoramas that are significantly better than other approaches. In addition, given the same input image, MultiConDiffusion can synthesize diverse and coherent results across multiple run, as is shown in Fig. 7.
Table 1 show the numerical comparison of MultiConDiffusion with other methods. We evaluate on 28 scenes provided by WonderJourney [36] and LucidDreamer [3]. Surprisingly, although our system does not incorporate image quality or aesthetics constraints, MultiConDiffusion can synthesis panorama with better quality and aesthetic score than other approaches. More importantly, panoramas generated by MultiConDiffusion show better consistency, aligning with our observation in Fig. 6 and demonstrating the effectiveness of our approach.
5.2 3D Scene Reconstruction Comparisons
Fig. 8 shows the visual comparison of PanoDreamer against LucidDreamer [3] and WonderJourney [36]. We set up their methods on the official code released by the authors, and use exactly the same training cameras as ours for a fair comparison. As is shown in the figure, neither LucidDreamer nor WonderJourney is capable of synthesizing coherent 360° 3D scene from a single input image. Their results suffer strong seam artifacts when the camera moves back to the beginning location. In contrast, PanoDreamer can synthesize consistent and seamless scene at any novel views.
Numerical results are shown in Table 2. The image quality and aesthetics of PanoDreamer results is slightly worse than WonderJourney and LucidDreamer, yet our consistency is significantly better. We argue that the rendering quality is more influenced by the consistency of the entire scene than the quality of each individual frame. For example, in a rendering video with 360° rotated camera, results of PanoDreamer appears to be coherent through the entire 360° scene without noticeable seam or content inconsistency, which is significantly better than other approaches. Please refer to our supplementary materials for video comparison and more visual results.
6 Conclusion

### v3

@@ 4.1. Single-Image Panorama Generation
During this stage, we fix $J_{T-1},\dots,J_{0}$ and find the optimal $L$ that minimizes Eq. 7. $L$ influences both the first term, as the diffusion model is conditioned on it (see Eq. 5), and the second term. Obtaining the optimal solution to each term independently is straightforward. The optimal solution to the first term is $L^{*}=L$, since if $L$ was used to produce the current $J_{T-1},\dots,J_{0}$, it is likely the best option for reproducing the same results. Moreover, since the second term is quadratic, the solution is simply $L^{*}=J_{0}$. Although obtaining the solution to each term is straightforward, computing the optimal solution considering both terms is difficult. However, assuming that $L$ and $J_{0}$ are close to each other—i.e., MultiConDiffusion does not diverge significantly from the condition image in one pass—it is reasonable to assume that $L^{*}=J_{0}$ is close to the optimal solution for both terms.
@@
We perform the optimization by first initializing $J_{T}$ with Gaussian noise and $L$ by placing the input image on a black canvas. We then alternate between stages 1 and 2 iteratively until convergence. At the end of this process, we can use either $\tilde{J}_{0}$ or $\tilde{L}$ as the final result. Fig. 3 compares MultiConDiffusion with MultiDiffusion and progressive inpainting.
Figure 4. We compare the result of our method, PanoDepthFusion, against applying Depth Anything V2 (DA V2) (Yang et al., 2024b) on the full image. The results obtained by DA V2 lacks details and is geometrically inconsistent. Our approach, on the other hand, produces highly detailed and consistent depth maps.
Figure 5. Averaging the patch depth estimates leads to banding artifacts since the depth maps are relative and not consistent. On the top right, we show that projecting the image into 3D using such a depth map results in clear banding artifacts. Since we initialize $G_{\theta_{i}}$ with the identity line, the patchwise average serves as our initial depth estimate during the optimization of Eq. 8. We also show our results after one and four iterations of optimization. After only four iterations, the seams disappear. As seen on the bottom right, the banding artifacts also disappear from the projected image.
Figure 6. We show the overview of PanoDepthFusion. We first apply an existing depth estimator to the overlapping patches of the input image to obtain a set of patch depth estimates. We then perform optimization in two stages. In the first stage, the depth patches are adjusted using a piecewise linear function $G_{\theta_{i}}$, and the adjusted patches are then aggregated to obtain the panoramic depth. In the second stage, we optimize the parameters $\theta_{i}$ of the parametric functions to match the adjusted patch depth estimates with the corresponding regions in the panoramic depth. These two steps are repeated until convergence.
Figure 7. We present an overview of our inpainting and 3DGS optimization process. Given a cylindrical panorama and its corresponding depth, we first convert them to the LDI representation. We then inpaint both the image and depth layers. Note that while all images and depth maps are cylindrical, we show only a small crop for clarity. Next, we initialize the Gaussians by assigning a single Gaussian to each pixel and projecting them into 3D space. Finally, we perform 3DGS optimization to obtain the 3D representation.
Table 1. Numerical comparison of MultiConDiffusion for single image wide-image generation against other relevant methods. CLIP-IQA+ and Q-Align measure the quality, A-CLIP and A-Align asses the aesthetic, and C-CLIP and C-Style evaluate the consistency of the results.
Method Q-IQA $\uparrow$ Q-Align $\uparrow$ A-CLIP $\uparrow$ A-Align $\uparrow$ C-CLIP $\uparrow$ C-Style $\downarrow$
Progressive 0.520 4.164 5.779 3.314 0.862 0.019
L-MAGIC (Cai et al., 2024) 0.550 4.331 5.865 3.426 0.842 0.069
MultiDiffusion (Bar-Tal et al., 2023) 0.523 4.390 5.953 3.516 0.869 0.030
SyncDiffusion (Lee et al., 2023) 0.535 4.290 5.893 3.429 0.864 0.016
MultiConDiffusion (ours) 0.530 4.481 5.992 3.696 0.881 0.011
4.1.1. Panorama Generation Details
@@
We slightly modify the MultiConDiffusion process to adapt it for generating panoramas from a single image. Our goal is to produce a cylindrical panorama, so in this case, MultiConDiffusion operates in the cylindrical domain, and the sequence $J_{T},\dots,J_{0}$ is defined within this domain. Since the base inpainting diffusion model operates on perspective images, $F_{i}$ performs both cropping and projection from the cylindrical to the perspective domain. Similarly, $F_{i}^{-1}$ projects the pixels from the perspective image back to the cylindrical image, placing them in the appropriate locations.
@@
We experimented with bilinear interpolation during the projection; however, interpolation smoothed out the noise, which negatively affected the performance of the diffusion model. Therefore, we instead use nearest-neighbor interpolation for both $F_{i}$ and $F_{i}^{-1}$. Additionally, we use an FOV of 45° for the perspective camera and carefully set the resolution of the cylindrical image to ensure a near one-to-one mapping between the pixels of the cylindrical and perspective images to preserve the noise pattern during projections.
@@
4.2. Panorama Depth Estimation
Given the panoramic image $\tilde{J}_{0}$, our goal is to estimate its depth $D$. In recent years, several powerful monocular depth estimation methods (Yang et al., 2024b, a) have been introduced. These approaches can estimate highly detailed relative depth but typically perform best at a specific image size. Beyond this optimal resolution, they often produce results that lack detail and geometric consistency. Consequently, applying these methods directly to panorama depth estimation leads to poor results, as shown in Fig. 4.
We address this problem by obtaining $D$ through a combination of estimated depth maps on patches using an existing technique, $\Psi$, i.e., $\Psi(F_{i}(\tilde{J}_{0}))$. However, naïvely combining the patches (e.g., through averaging) leads to unsatisfactory results (see Fig. 5), as the patch depth estimates are relative and can be inconsistent. To overcome this, we pose the problem of obtaining panoramic depth from patch depth estimates as an optimization task. Our key insight is that the panoramic depth $D$ should be close to the estimated depth after it has been globally aligned through a parametric function. This can be formally written as:
(8) $D^{*},\bm{\theta}^{*}=\operatorname*{arg\,min}_{D,\bm{\theta}}\sum_{i=1}^{n}\|F_{i}(D)-G_{\theta_{i}}(\Psi(F_{i}(\tilde{J}_{0})))\|^{2},$}
@@ 4.2. Panorama Depth Estimation
Solving for both $D$ and $\bm{\theta}$ simultaneously is challenging. Therefore, we propose performing this optimization through alternating minimization, consisting of two stages (see Fig. 6). In the first stage, we fix $\bm{\theta}$ and find the optimal $D$. Since the objective is quadratic, the solution can be obtained in closed form, similar to Eq. 3. The only difference is that $\Phi$, the diffusion model, is replaced with $\Psi$, and $W_{i}=\bm{1}$. In the second stage, we fix $D$ and find the optimal $\bm{\theta}$. This is a least-squares regression problem, which we solve using standard packages.
@@
4.3. Inpainting and 3DGS Optimization
@@
We now discuss our process for reconstructing the 3D scene using our generated panorama and the corresponding depth map (see the overview in Fig. 7). While the estimated depth can be used to project the cylindrical image into 3D space, when the scene is viewed from any position other than the panorama’s center of projection, occluded regions become visible. To address this, we utilize the layered depth image (LDI) inpainting approach by Shih et al. (2020), which performs effective depth-aware texture inpainting while also providing the corresponding depth. We use a four-layer LDI representation (foreground, background, and two intermediate layers) based on agglomerative clustering by disparity.
@@
We then use these cylindrical layered images and depth maps to initialize a set of 3D Gaussians. Specifically, we assign a Gaussian to each pixel of the image at each layer and project them into 3D according to the corresponding depth. The color of each Gaussian is initialized based on the pixel color (without spherical harmonics); we initialize the rotation matrix with identity, assign the scale following Paliwal et al. (2025), and set the opacity to 0.5. During this process, we keep track of which Gaussians correspond to which layer, as this information is required for optimization.
@@
Next, we perform 3DGS optimization for 1000 iterations. To do this, we set up 240 evenly rotated cameras from the center of projection and project the layered images and depth maps to these cameras. During the optimization, we randomly sample one of these cameras and optimize the Gaussians according to their corresponding layer. Additionally, we composite all the four layers and use the composited image as a reference to optimize all the Gaussians. Note that per layer and composite losses are optimized simultaneously. We use the original 3DGS reconstruction loss along with an $L_{2}$ loss between the rendered and layered depth maps. In addition, to be able to produce consistent results from novel views, we use the depth-based novel view loss, proposed by Zhu et al. (2024). Furthermore, pruning and densification are performed following 3DGS. Once the optimized 3DGS representation is obtained, we can synthesize novel views of the scene and produce coherent and seamless results.
Figure 8. We compare the wide-images generated by MultiConDiffusion with those from other methods. Other approaches often result in sharp discontinuities and contextual inconsistencies. For instance, in the top example, the MultiDiffusion result shows a mismatch between the generated sky and the input sky.
5. Results
@@
In this section, we evaluate both the generated texture and depth of our proposed algorithms. Specifically, we compare our approach against state-of-the-art wide-image generation and 3D scene generation methods, both visually and numerically. For evaluation, we compile a test set of 28 real and synthetic scenes sourced from LucidDreamer (Chung et al., 2023) and WonderJourney (Yu et al., 2024b). For numerical evaluation, we employ several metrics to evaluate different aspects of the results: (1) Quality - we assess the quality of results using CLIP-IQA+ (Wang et al., 2023) and Q-Align (Wu et al., 2023a) scores. CLIP-IQA+ and Q-Align are built upon contrastive language-image pre-training (CLIP) (Radford et al., 2021) and large multi-modality models (LMMs) for image quality assessment, respectively. (2) Aesthetic - we use the CLIP aesthetic score (A-CLIP) and A-Align (Wang et al., 2023) to measure results aesthetic. (3) Consistency - we compute the similarity (C-CLIP) and style loss (Gatys et al., 2016) (C-Style) of the CLIP embeddings of random pairs of non-overlapping patches in the results.
Table 2. Numerical comparisons of our approach against the state-of-the-art methods on novel view synthesis. The evaluation metrics are the same as the ones in Table 1.
Method Q-IQA $\uparrow$ Q-Align $\uparrow$ A-CLIP $\uparrow$ A-Align $\uparrow$ C-CLIP $\uparrow$ C-Style $\downarrow$
LucidDreamer (Chung et al., 2023) 0.495 2.911 5.253 2.705 0.848 0.058
WonderJourney (Yu et al., 2024b) 0.504 3.506 5.368 2.834 0.820 0.058
PanoDreamer (ours) 0.443 3.305 5.673 2.772 0.869 0.025
Figure 9. We compare renderings of PanoDreamer with LucidDreamer (Chung et al., 2023) and WonderJourney (Yu et al., 2024b). For each methods, we render 3D scene from two novel views. LucidDreamer and WonderJourney produces results with seams and inconsistencies. In comparison, PanoDreamer is capable of generating coherent renderings from novel views. For more visual results and video comparison, please refer to our supplementary materials.
Table 3. Comparison of PanoDepthFusion with Moge’s (Wang et al., 2025) patch-wise depth estimation + blending approach. Rows 232-3 correspond to PanoDepthFusion combined with different depth estimators.
Method AbsRel $\downarrow$ RMSE $\downarrow$ $\delta_{1}\uparrow$ $\delta_{2}\uparrow$ $\delta_{3}\uparrow$
MoGe 0.21 0.56 61.66 90.63 98.32
DA V2 + ours 0.31 0.91 47.05 77.39 92.17
MoGe + ours 0.19 0.67 66.47 92.66 98.71
Table 4. Numerical ablations to evaluate the effectiveness of different components of our novel view synthesis pipeline. The evaluation metrics are the same as the ones in Table 1.
Method Q-IQA $\uparrow$ Q-Align $\uparrow$ A-CLIP $\uparrow$ A-Align $\uparrow$ C-CLIP $\uparrow$ C-Style $\downarrow$
LucidDreamer (Chung et al., 2023) 0.495 2.911 5.253 2.705 0.848 0.058
LuciDreamer w/ our panorama 0.469 3.012 5.465 2.871 0.862 0.022
Ours + LDI 0.469 2.546 5.479 2.542 0.863 0.016
PanoDreamer (ours w/ 3DGS) 0.443 3.305 5.673 2.772 0.869 0.025
Moreover, we numerically evaluate our depth estimation against other techniques, on 30 randomly selected scenes from the real-world Stanford2D3D dataset (Armeni et al., 2017). For this evaluation, we use Absolute Relative Error (AbsRel), Root Mean Squared Error (RMSE), and three percentage metrics, i.e., the percentages of pixels where the ratio ($\delta$) between the estimated and ground truth depth is smaller than $1.25$, $1.25^{2}$, and $1.25^{3}$.
5.1. Wide-Image Reconstruction Comparisons
Table 1 numerically compares MultiConDiffusion against vanilla progressive inpainting using our base inpainting diffusion model, L-MAGIC (Cai et al., 2024), SyncDiffusion (Shih et al., 2020), and MultiDiffusion (Bar-Tal et al., 2023). Note that, since these images are not as wide as cylindrical panoramas, we perform the optimization for 15 iterations instead of 20. As seen, our method produces better results than all the other approaches across nearly all metrics. More importantly, the images generated by MultiConDiffusion show better consistency in terms of both style and content, demonstrating the effectiveness of our optimization strategy.
In Fig. 8, we show visual comparison against the other approaches. As seen, progressive inpainting generates results with noticeable seams and strong inconsistency. L-Magic which works based on progressive inpainting, gradually changes the style of the image closer to the two sides. Similarly SyncDiffusion and MultiDiffusion produce results that are not consistent with the input images. Note that the walkway in center of Multidiffusion’s result does not align with the surrounding regions. In contrast, MultiConDiffusion can generate coherent and seamless wide-images that are significantly better than other approaches.
5.2. Panorama Depth Estimation
In Table 3, we numerically evaluate the performance of PanoDepthFusion with two baseline depth estimation methods: Depth Anything V2 (Yang et al., 2024b) (“DA V2 + ours”) and MoGe (Wang et al., 2025) (“MoGe + ours”), a recently introduced (concurrent) single-image depth estimation method. As seen, our approach with MoGe produces better results than DA V2. We also report the metrics for MoGe with its own patch-wise depth estimation and blending approach (“MoGe”). Comparing “MoGe” with “MoGe + ours” demonstrates that PanoDepthFusion achieves better performance across most metrics. For our 3D scene reconstruction experiments, we adopt DA V2 as the base estimator, since it performs reliably across diverse settings (indoor, outdoor, synthetic) and integrates well with our pipeline. Nonetheless, our approach is agnostic to the choice of depth estimator and can be combined with MoGe or future methods, potentially yielding further improvements.
5.3. 3D Scene Reconstruction Comparisons
We show numerical comparisons of our PanoDreamer against LucidDreamer (Chung et al., 2023) and WonderJourney (Yu et al., 2024b) in Table 2. We use the official code released by the authors, and utilize the same training cameras as ours for a fair comparison. While our image quality and aesthetic scores are slightly worse than WonderJourney, our consistency scores are significantly better. This is because their novel view images are often reasonable when viewed one image at a time, however, different novel view images differ in style and thus are not consistent. Our approach, on the other hand, produces results that are consistent across all views.
We further show visual comparisons against the other methods in Fig. 9. LucidDreamer and WonderJourney produce results with seams and inconsistent style across the two views shown here. In contrast, PanoDreamer can synthesize consistent and seamless scene at both novel views. Please refer to our supplementary materials for video comparison and more visual results.
5.4. Novel View Synthesis Ablations
We evaluate the effectiveness of different components in our 3D reconstruction pipeline on novel view synthesis task, both visually (Fig 10) and numerically (Table 4). While incorporating our coherent panorama improves LucidDreamer’s performance, their generated novel views still contain artifacts due to inconsistent depth estimation. Moreover, we show the impact of 3DGS optimization by comparing our results against LDI rendering. As seen in Fig. 10, LDI’s rendering contains artifacts around the depth layer discontinuities. These are significantly reduced with our 3DGS optimization partly due to our use of the depth-based novel view loss.
Figure 10. We highlight the effect of various components of our novel view synthesis pipeline. Using our consistent panorama in progressive techniques like LucidDreamer (Chung et al., 2023) improves the texture quality around input image boundaries. However, the rendered texture still contains artifacts due to inconsistent scene depth estimation. Moreover, using the intermediate layered depth image (LDI) representation instead of 3DGS (Ours) leads to subpar rendered images with with artifacts around the depth layer discontinuities.
6. Conclusion, Limitation, and Future Work